Calculate the volume (in mL) of anisaldehyde that is the molar equivalent of 0.20 mL of benzaldehyde.
Answer #1

The density of benzaldehyde = 1.04 g/mL

Mass/Volume = Density, Mass = Volume X Density

Mass of benzaldehyde = 0.20 mL X 1.04 g/mL = 0.21 grams

Molar mass of Benzaldehyde = 106.12 g/mol

Moles of benzaldehyde = Mass/Molar mass = 0.21/106.12 = 0.00196 moles

Similarly as earlier, Mass of anisaldehyde = Volume(in mL) X 1.12 g/mL

Molar mass of anisaldehyde = 136.15 g/mol

Moles of anisaldehyde = (Volume X 1.12)/136.15 moles

As both need to be molar equivalent, Moles of benzaldehyde = Moles of anisaldehyde

0.00196 = (Volume X 1.12)/136.15

Volume of anisaldehyde = (0.00196 X 136.15)/1.12

Volume of anisaldehyde = 0.24 mL